OpenStudy (anonymous):

solve logarithmic function algebraically: ln(x+5) = ln(x-1) - ln(x+1)

Nnesha (nnesha):

\[\log a - \log b = \log \frac{ a }{ b}\] subtraction change to division addition change to multiplication ^^these are soe rules

Nnesha (nnesha):

so who would you change right side ????^

OpenStudy (anonymous):

What would you do with the ln(x+5)?

Nnesha (nnesha):

fist we need to solve right change that into division then you cancel out ln both side

OpenStudy (anonymous):

What am I canceling out?

Nnesha (nnesha):

 you got this on right side ??